# Advancing the Database of Cross-Linguistic Colexifications with New Workflows and Data

### v1

@@ 3.3 Language and Concept Selection
After detecting this problem when reviewing individual datasets in CLICS 3, we decided to modify the criterion for the selection of languages in three ways. First, instead of setting the threshold to 250 words per language, we lowered it to 180 words, accounting for the fact that almost half of the languages in CLICS 3 would not pass this threshold. Second, in our modified data aggregation workflow, the threshold is applied to individual language varieties rather than to entire datasets. This means that for all languages in the sample, we count if they meet the inclusion criterion or not. As a result, it may happen that only certain parts of the datasets from which CLICS* aggregates the word lists make it into the final database. Third, in order to yield a more meaningful selection of concepts, our workflow first orders all concepts by their occurrence across the languages in the data and then retains the most frequent 1,800 concepts. When aggregating the data from the individual word lists, only these concepts are retained. This procedure helps to decrease the scarcity of the data, resulting from the fact that the individual word lists often differ quite drastically with respect to the concepts for which they provide translations.
@@ 3.4 Data Representation
When following the core principle of CLDF in using tables as the basic representation format wherever possible, it would be straightforward to represent a graph with the help of two tables. One table would represent the nodes of a graph, with node attributes being provided in additional columns, and another table would represent the edges, with edge attributes being represented in additional columns. It turned out that this format could not only be easily represented in the CLDF specification, but that it would allow us to represent colexification data in the form of a structural dataset in which certain typological properties are provided as structural information for individual languages. Following Forkel and List (2020) in combining a word list and a structural dataset in a unified CLDF dataset, CLICS* now consists of a large aggregated word list with individual word forms across several thousand language varieties along with structural data that provides information on the languages that exhibit certain colexifications.
@@
Structural data in CLDF typically consist of a parameter table that provides information on the features comparable across languages, and a value table that provides information on the individual values as they are reflected in individual languages. In our new data model for cross-linguistic colexification data, all individual colexifications that can be inferred when analyzing the aggregated word list feature as parameters, whereas the corresponding values for each language are represented by three different codes, indicating if the feature represented by the parameter is present, absent, or missing. Thus, our proposal for CLICS* not only informs whether a given language exhibits a particular colexification but also whether it does not show the colexification, or whether the information is missing, since translations for at least one of the concepts involved in the colexification are missing in the word list.
@@
There are two major advantages of this new representation. The first advantage is that colexifications can be directly inspected in tabular form. Since the colexification data are shared in a table format as part of the CLDF dataset underlying CLICS*, interested users can browse through the colexifications using their favorite spreadsheet editor. Analyzing the colexification network with software tools is also facilitated, given that all major tools support tabular data. This means that networks can be conveniently analyzed computationally or visualized with graph visualization software, such as Cytoscape (see, for example, the tutorial by Tjuka, 2024a). The second advantage is that it is much easier to integrate the data produced by CLICS* with the data shared by other projects. Community assignments along with additional information on the coverage of concepts across languages and language families, for example, are now part of the concept table that serves as the basic parameter table for the CLICS* word list. From this representation, it is easy to integrate the data not only into the Concepticon (see also Bocklage et al., 2024) but also into extended reference catalogs such as NoRaRe (https://norare.clld.org, Tjuka et al., 2022), a catalog that extends the Concepticon by providing additional information on norms, rates, and ratings for words and concepts across multiple languages.
@@ 3.5 Implementation
CLICS* is implemented in the form of a CLDFBench package (Forkel and List, 2020), written in Python, that can be installed from the command line and contains the resulting CLDF data along with the code that was used to create the data. The package is shared as part of the supplemental material accompanying this study and contains additional information and code examples that were used to produce the findings presented in this study.
Figure 2: CLICS* colexification network with two selected communities (central concepts DRESS and EDGE).
@@ 4 Data Validation
4.1 Comparing CLICS 3 and CLICS*
In order to understand the differences between our updated version CLICS* and the previous versions of CLICS, most importantly the last officially published version CLICS 3 by Rzymski et al. (2020), we carried out a detailed comparison of CLICS 3 and CLICS*. Given that we deliberately restricted the number of concepts in CLICS* to an initial list of 1,800 concepts – of which 1,730 were retained when selecting those languages that would cover at least 180 concepts of the initial list – it may seem as if CLICS* simply reduced the amount of data in contrast to CLICS 3. However, this is not quite the case, which is apparent when comparing the number of words, language varieties, languages (different glottocodes), and language families covered in both datasets, as shown in Table 1. CLICS* exceeds CLICS 3 not only regarding the number of language families and language varieties covered, but most notably with respect to the number of word forms that are provided in phonetic transcriptions. CLICS* reaches almost the same size as CLICS 3, while providing almost three times as many phonetics transcriptions.
A similar situation arises when comparing the overall number of concepts with the average number of languages and families expressing a concept in both datasets (also shown in Table 1). Here, CLICS 3 exceeds CLICS* in the number of concepts that are colexified (1,647 vs. 1,386), while showing similar values for the average number of languages expressing a concept (624 vs. 607). However, regarding the average number of families expressing a concept, CLICS* largely exceeds CLICS 3 (92 vs. 61).
In sum, the comparison provided in Table 1 shows that CLICS* does not simply provide more data, resulting in more languages, more concepts, and more colexifications. Instead, the major improvements concerning the data basis, concept handling, and language selection yield a colexification network that consolidates the tendencies in the data rather than to diversify them further. Thus, while CLICS* has fewer colexified concepts, the concepts in the colexification network of CLICS* have more connections across more language families on average, as reflected in their degree distribution (6 vs. 5), and these connections are also substantiated by more examples, as reflected in the weighted degree distribution (53 vs. 36). This trend can also be observed when comparing the inferred colexifications directly. There are 2,874 colexifications observed in both networks, 1,354 unique to CLICS 3, and 1,111 unique to CLICS*. Of the common edges, 859 colexifications in CLICS* can be found in more language families, compared to 778 colexifications in CLICS 3.
Criterion CLICS 3 CLICS*
Datasets 30 94
Varieties 3 156 3 420
Languages 2 280 2 150
Families 200 249
Words 1 462 125 1 443 325
Transcriptions 563 878 1 443 325
Words per Variety 467 422
@@ 4.1 Comparing CLICS 3 and CLICS*
Table 1: Comparison between CLICS 3 and CLICS*. Colexifications are only counted when occurring in at least three different language families. Weighted degree is calculated by counting the number of language families per link.
4.2 Visualizing the CLICS* Network
To create a visual representation of the CLICS* network, researchers can either use the GML file that is provided along with the CLDF data, or the table with all colexifications that is shared as part of CLDF directly. As mentioned in § 3.4, the new data representation in tabular form as part of a unified CLDF dataset makes it easier to analyze the data computationally. The visualization of the data is also greatly facilitated, given that edge tables are the basic input format for network visualization software tools like Cytoscape. A tutorial on how to create a network visualization with Cytoscape is provided in Tjuka 2024a. We used this approach to create Figure 2, which provides a bird’s eye view of the CLICS* network.
The figure shows the entire network with two communities highlighted and enlarged. The first community has the concept dress as a central node and shows colexifications with other clothing items. The edge weights represent the frequency in which a given colexification occurs across languages. For example, the colexification between dress and skirt is more frequent than the colexification between dress coat. The second community has the concept edge as a central node and includes cross-linguistically frequent colexifications such as edge and side and less frequent ones like edge and side. Given the straightforward representation of the colexification network in CLICS*, the data can conveniently be explored. By using Cytoscape, researchers can further investigate the properties of the network and filter them according to their particular research interests.
@@ 5 Conclusion and Outlook
We presented CLICS*, an enhanced version of the Cross-Linguistic Colexification Database, which integrates lexical data for 3,420 language varieties, corresponding to 2,150 distinct Glottocodes. When creating CLICS* we used an advanced workflow for the aggregation and analysis of cross-linguistic colexification data that is based on an increased and improved data basis, an improved handling of concepts, more fine-grained criteria for the selection of languages and concepts, and an updated representation of the colexification data.
In contrast to previous colexification databases, CLICS* determines colexifications exclusively based on phonetic transcriptions. This makes the data more consistent and robust and opens new possibilities to analyze the data in comparative studies. Due to the phonetic transcriptions in CLICS*, future studies can build on the initial work to infer and investigate partial colexifications (List, 2023; Tjuka and List, 2024; Rubehn and List, 2025). In addition, phonetic transcriptions enable scholars to carry out more fine-grained analyses of colexifications inside specific language families, where a handling of cognate words is important to identify colexifications that have evolved independently from colexifications that have been inherited across branches (Tjuka et al., 2024).
Future studies can use CLICS* to explore the relationship between words and their meanings across a wide range of languages and uncover important insights into language evolution, cultural variations, and cognitive principles. In this way, CLICS* has great potential to contribute to future studies that address open questions in a broad range of linguistic subfields, including linguistic typology, historical linguistics, psycholinguistics, and computational linguistics.
@@
All data and code underlying this study along with instructions on how to run the code are available on Open Science Framework. They can be downloaded here: https://osf.io/8gqtn/?view_only=8be8f04428d34409 ad59cf00deccfaa2.
@@
General limitations that apply to large-scale aggregation studies in comparative linguistics apply also to CLICS*. These include the fact that the word list approach for aggregating cross-linguistic colexifications may fail to model fine-grained aspects of colexifications in individual language families, many of which cannot be modeled appropriately without a detailed inspection of particular languages and their history. An additional problem of all cross-linguistic colexification databases is that they contain a lot of missing data, showing low coverage for most concepts cross-linguistically. We also emphasize that detailed studies investigating the properties of CLICS* are missing so far, but we envisage that these will be carried out by different teams (not only including the team which compiled the data by now).
@@ Acknowledgments
This project was supported by the ERC Consolidator Grant ProduSemy (JML; Grant No. 101044282, see https://doi.org/10.3030/101044282) and the ERC Synergy Grant QUANTA (CR; Grant No. 951388, see https://doi.org/10.3030/951388). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency (nor any other funding agencies involved). Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ 3.3 Language and Concept Selection
After detecting this problem when reviewing individual datasets in CLICS 3, we decided to modify the criterion for the selection of languages in three ways. First, instead of setting the threshold to 250 words per language, we lowered it to 180 words, accounting for the fact that almost half of the languages in CLICS 3 would not pass this threshold. The threshold was chosen because we noticed that there were many datasets with 200 words or fewer. For many languages, only versions of the Swadesh list with 200 concepts (Swadesh, 1952) are available, so the chance of obtaining some concepts missing for individual languages is considerably high. Setting the threshold a bit lower allows us to predefine a core set of concepts that are comparable across languages (and which could be modified anytime, depending on the analysis one desires to conduct). Second, in our modified data aggregation workflow, the threshold is applied to individual language varieties rather than to entire datasets. This means that for all languages in the sample, we count whether they meet the inclusion criterion or not. As a result, it may happen that only certain parts of the datasets from which CLICS 4 aggregates the word lists make it into the final database. Third, in order to yield a more meaningful selection of concepts, our workflow first orders all concepts by their occurrence across the languages in the data and then retains the most frequent 1,800 concepts. When aggregating the data from the individual word lists, only these concepts are retained. This procedure helps to decrease the sparsity of the data, resulting from the fact that the individual word lists often differ quite drastically with respect to the concepts for which they provide elicitation glosses. While the cutoff point may seem arbitrary, it reflects our experience in working with the mapping of concept lists in the Concepticon project: beyond 1,800 concepts, the chances of finding concepts expressed across many languages from many different families drop considerably.
@@ 3.4 Data Representation
When following the core principle of CLDF in using tables as the basic representation format wherever possible, it would be straightforward to represent a graph with the help of two tables. One table would represent the nodes of a graph, with node attributes being provided in additional columns, and another table would represent the edges, with edge attributes being represented in additional columns. It turned out that this format could not only be easily represented in the CLDF specification, but that it would allow us to represent colexification data in the form of a structural dataset (Forkel et al., 2018). While the primary dataset underlying CLICS 4 provides information on colexifications between a fixed set of standardized concept sets, the additional view as a structural dataset – resembling a cross-linguistic typological database – offers a language-centered view: colexifications are modeled as parameters and for each language we provide information on their presence or absence. Thus, following Forkel and List (2020) in combining a word list and a structural dataset in a unified CLDF dataset, CLICS 4 now consists of a large aggregated word list with individual word forms across several thousand language varieties, along with structural data that provides information on the languages that exhibit certain colexifications.
@@
Structural data in CLDF typically consist of a parameter table that provides information on the features comparable across languages, and a value table that provides information on the individual values as they are reflected in individual languages. In our new data model for cross-linguistic colexification data, all individual colexifications that can be inferred when analyzing the aggregated word list feature are represented as parameters. In contrast, the corresponding values for each language are represented by three different codes, indicating if the feature represented by the parameter is present, absent, or missing. Thus, our proposal for CLICS 4 not only informs whether a given language exhibits a particular colexification but also whether it does not show the colexification, or whether the information is missing, since elicitation glosses for at least one of the concepts involved in the colexification are missing in the word list.
@@
There are two major advantages of this new representation. The first advantage is that colexifications can be directly inspected in tabular form. Since the colexification data are shared in a table format as part of the CLDF dataset underlying CLICS 4, interested users can browse through the colexifications using their favorite spreadsheet editor. Analyzing the colexification network with software tools is also facilitated, given that all major tools support tabular data. This means that networks can be conveniently analyzed computationally or visualized with graph visualization software, such as Cytoscape (for a tutorial, see Tjuka, 2024a). The second advantage is that it is much easier to integrate the data produced by CLICS 4 with the data shared by other projects. Community assignments, along with additional information on the coverage of concepts across languages and language families, for example, are now part of the concept table that serves as the basic parameter table for the CLICS 4 word list. From this representation, it is easy to integrate the data not only into the Concepticon (see also Bocklage et al., 2024) but also into extended reference catalogs such as NoRaRe (https://norare.clld.org, Tjuka et al., 2022), a catalog that extends the Concepticon by providing additional information on norms, rates, and ratings for words and concepts across multiple languages.
@@ 3.5 Implementation
CLICS 4 is implemented in the form of a CLDFBench package (Forkel and List, 2020), written in Python, that can be installed from the command line and contains the resulting CLDF data along with the code that was used to create the data. The package is shared as part of the supplemental material accompanying this study and contains additional information and code examples that were used to produce the findings presented in this study.
Figure 2: CLICS 4 colexification network with two selected communities (central concepts DRESS and EDGE).
@@ 4 Data Validation
4.1 Comparing CLICS 3 and CLICS 4
In order to understand the differences between our updated version CLICS 4 and the previous versions of CLICS, most importantly the last officially published version CLICS 3 by Rzymski et al. (2020), we carried out a detailed comparison of CLICS 3 and CLICS 4. Given that we deliberately restricted the number of concepts in CLICS 4 to an initial list of 1,800 concepts – of which 1,730 were retained when selecting those languages that would cover at least 180 concepts of the initial list – it may seem as if CLICS 4 simply reduced the amount of data in contrast to CLICS 3. However, this is not the case, which is apparent when comparing the number of words, language varieties, languages (different glottocodes), and language families covered in both datasets, as shown in Table 1. CLICS 4 exceeds CLICS 3 not only regarding the number of language families and language varieties covered, but most notably with respect to the number of word forms that are provided in phonetic transcriptions. CLICS 4 reaches almost the same size as CLICS 3, while providing almost three times as many phonetic transcriptions.
A similar situation arises when comparing the overall number of concepts with the average number of languages and families expressing a concept in both datasets (also shown in Table 1). Here, CLICS 3 exceeds CLICS 4 in the number of concepts that are colexified (1,386 vs. 1,647), while showing similar values for the average number of languages expressing a concept (607 vs. 624). However, regarding the average number of families expressing a concept, CLICS 4 largely exceeds CLICS 3 (92 vs. 61).
In sum, the comparison provided in Table 1 shows that CLICS 4 does not simply provide more data, resulting in more languages, more concepts, and more colexifications. Instead, the major improvements concerning the data basis, concept handling, and language selection yield a colexification network that consolidates the tendencies in the data rather than diversifying them further. Thus, while CLICS 4 has fewer colexified concepts, i.e., concepts that are part of a colexification, the concepts in the colexification network of CLICS 4 have more connections across more language families on average, as reflected in their degree distribution (6 vs. 5). In addition, these connections are also substantiated by more colexifications, as reflected in the weighted degree distribution (53 vs. 36). This trend can also be observed when directly comparing the inferred colexifications. There are 2,874 colexifications observed in both networks, 1,354 unique to CLICS 3, and 1,112 unique to CLICS 4. Of the common edges, 859 colexifications in CLICS 4 can be found in more language families, compared to 778 colexifications in CLICS 3.
Criterion CLICS 3 CLICS 4
Datasets 30 95
Varieties 3 156 3 432
Languages 2 280 2 152
Families 200 247
Words 1 462 125 1 445 845
Transcriptions 563 878 1 445 845
Words per Variety 467 421
@@ 4.1 Comparing CLICS 3 and CLICS 4
Table 1: Comparison between CLICS 3 and CLICS 4. Colexifications are only counted when occurring in at least three different language families. Weighted degree is calculated by counting the number of language families per link.
4.2 Visualizing the CLICS 4 Network
To create a visual representation of the CLICS 4 network, researchers can either use the GML file that is provided along with the CLDF data, or the table with all colexifications that is shared as part of CLDF directly. As mentioned in § 3.4, the new data representation in tabular form as part of a unified CLDF dataset makes it easier to analyze the data computationally. The visualization of the data is also greatly facilitated, given that edge tables are the basic input format for network visualization software tools like Cytoscape. A tutorial on how to create a network visualization with Cytoscape is provided in Tjuka 2024a. We used this approach to create Figure 2, which provides a bird’s eye view of the CLICS 4 network.
The figure shows the entire network with two communities highlighted and enlarged. The first community has the concept dress as a central node and shows colexifications with other clothing items. The edge weights represent the frequency with which a given colexification occurs across languages. For example, the colexification between dress and skirt is more frequent than the colexification with coat. The second community has the concept edge as a central node and includes cross-linguistically frequent colexifications such as edge and side and less frequent ones like edge and corner. Given the straightforward representation of the colexification network in CLICS 4, the data can conveniently be explored. By using Cytoscape, researchers can further investigate the properties of the network and filter them according to their particular research interests.
@@ 5 Conclusion and Outlook
We presented CLICS 4, an enhanced version of the Cross-Linguistic Colexification Database, which integrates lexical data for 3,432 language varieties, corresponding to 2,152 distinct Glottocodes. When creating CLICS 4 we used an advanced workflow for the aggregation and analysis of cross-linguistic colexification data that is based on an increased and improved data basis, an improved handling of concepts, more fine-grained criteria for the selection of languages and concepts, and an updated representation of the colexification data.
In contrast to previous colexification databases, CLICS 4 determines colexifications exclusively based on phonetic transcriptions. This makes the data more consistent and robust and opens new possibilities to analyze the data in comparative studies. Due to the phonetic transcriptions in CLICS 4, future studies can build on the initial work to infer and investigate partial colexifications (List, 2023; Tjuka and List, 2024; Rubehn and List, 2025). In addition, phonetic transcriptions enable scholars to carry out more fine-grained analyses of colexifications inside specific language families, where a handling of cognate words is important to identify colexifications that have evolved independently from colexifications that have been inherited across branches (Tjuka et al., 2024).
Future studies can use CLICS 4 to explore the relationship between words and their meanings across a wide range of languages and uncover important insights into language evolution, cultural variations, and cognitive principles. In this way, CLICS 4 has great potential to contribute to future studies that address open questions in a broad range of linguistic subfields, including linguistic typology, historical linguistics, psycholinguistics, and computational linguistics.
@@
All data and code underlying this study, along with instructions on how to run the code, are openly available. The CLICS 4 database is curated on GitHub (https://github.com/clics/clics4/tree/v0.5, Version 0.5) and archived with Zenodo (DOI: https://doi.org/10.5281/zenodo.16900180). The code that we used to compare CLICS 3 and CLICS 4 is cuarated on Codeberg (https://codeberg.org/calc/clics4-paper/src/tag/v1.0, Version 1.0) and archived with Zenodo (DOI: https://doi.org/10.5281/zenodo.16902185).
@@
General limitations that apply to large-scale aggregation studies in comparative linguistics also apply to CLICS 4. These include the fact that the word list approach for aggregating cross-linguistic colexifications may fail to model fine-grained aspects of colexifications in individual language families, many of which cannot be modeled appropriately without a detailed inspection of particular languages and their history. An additional problem of all cross-linguistic colexification databases is that they contain a lot of missing data, showing low coverage for most concepts cross-linguistically. We also emphasize that detailed studies investigating the properties of CLICS 4 are missing so far, but we envisage that these will be carried out by different teams (not only including the team which compiled the data by now). Another improvement that needs to be implemented in the future is the treatment of some artificially separated concepts. For example, the current version splits the concept think into the more specific concepts think (reflect) and think (believe). While this modification reflects the ambiguity of the concept think, we suspect that there is no frequently used questionnaire for cross-linguistic data that would contain both think (reflect) and think (believe). As a result, one may call the colexification between think (reflect) and think (believe) in question, given that the database lacks direct evidence. This holds to an even larger degree for kinship terms. One solution we could think of would be to consider only think, as the broadest concept, because this concept is present in most languages. While our current technology would allow for such a handling, we think addressing this problem in a principled way will require a more thorough revision, potentially accompanied by additional computational analyses and very detailed decisions that should not be made in an ad-hoc style.
So far, CLICS 4 is limited to the data and the database itself can only be investigated with tools for network visualization and with computational approaches. As of now, the web application at https://clics.clld.org still serves the data underlying CLICS 3. Implementing the web application for CLICS 4 is planned and will follow in the near future.
@@ Acknowledgments
This project was supported by the ERC Consolidator Grant ProduSemy (JML; Grant No. 101044282, see https://doi.org/10.3030/101044282) and the ERC Synergy Grant QUANTA (CR; Grant No. 951388, see https://doi.org/10.3030/951388). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency (nor any other funding agencies involved). Neither the European Union nor the granting authority can be held responsible for them.